OpenStudy (anonymous):

describe the domain of the following function f(x)=-6x+4 f(x)=sqrt10-x

OpenStudy (anonymous):

Domain of a line is all real numbers, unless its vertical

OpenStudy (anonymous):

can you tell that just by looking at it?

OpenStudy (anonymous):

Yes f(x)=mx+b is the form of a line, x=a is the form of a vertical line. So for the first one its all real numbers Think of it like this, you can plug any integer, positive or negative into -6x+4 and get an integer out

OpenStudy (anonymous):

because even if x = 0 then you still get 4?

OpenStudy (anonymous):

Yep you can plug in all real numbers and you will get a real number output. for the second one is it \[\sqrt{10-x}\]

OpenStudy (anonymous):

what does "no domain restrictions" mean?

OpenStudy (anonymous):

What i just said, you can plug in all real numbers and you will get a real number output

OpenStudy (anonymous):

oh ok. gotcha. thanks

OpenStudy (anonymous):

For the second one, under the square root has to be greater than or equal to 0 so to find the domain solve this little inequality \[10-x \ge0\]

OpenStudy (anonymous):

x=10, so it is \[\ge 0\]

OpenStudy (anonymous):

or actually I didnt do that right.

OpenStudy (anonymous):

in fact I did it quite opposite didn't I?

OpenStudy (anonymous):

x=-10

OpenStudy (anonymous):

maybe I'm getting ahead of myself. I forgot that I divided out a -1 from the -10 which is x=10

OpenStudy (anonymous):

Its an inequality not an equation so first subtract 10 from both sides to get \[-x \ge -10\]

OpenStudy (anonymous):

then divide by -1, which flips the inequality sign \[x \le 10\]

OpenStudy (anonymous):

oh, so when you divide the neg 1 (or any number?) from the x to the 10 you flip the sign. I always forget that rule

OpenStudy (anonymous):

Yeah just gotta burn it in your brain. When you divide or multiply by a negative number it flips the inequality so your domain is all x less than or equal to 10

OpenStudy (anonymous):

or ( -\[\infty \],10]

OpenStudy (anonymous):

that didn't format right, but you get me.

OpenStudy (anonymous):

Yes thats correct \[(-\infty,10]\]

OpenStudy (anonymous):

what if it's g(x)= \[\frac{ 1 }{ x^{2}+4x-5 }\]

OpenStudy (anonymous):

You cant divide by 0, so set the denominator equal to 0 and solve. What you get will be the values that arent in the domain

OpenStudy (anonymous):

ok, after I get it to x^2+4x-5 I do the quadratic formula?

OpenStudy (anonymous):

or, complete the square I mean...

OpenStudy (anonymous):

Use whatever method you like best. It factors easy though (x-1)(x+5)=0

OpenStudy (anonymous):

gotcha, ok. cool

OpenStudy (anonymous):

you wouldn't happen to know of a good site with a listing of all these rules for the domains would you?

OpenStudy (anonymous):

if not, I think I have them in my notes, I'll find them and look over it. thanks for your help!

OpenStudy (anonymous):

Hmm, you could try googling it. But basically, if theres a square root involved, you know under the square root has to be 0 or positive. if there's a fraction you know the denominator cant be 0, things like that

OpenStudy (anonymous):

ok, thanks again